# Extracting Conceptual Knowledge to Locate Software Issues

### v1

@@ Abstract.
Issue localization, which identifies faulty code elements such as files or functions, is critical for effective bug fixing. While recent LLM-based and LLM-agent-based approaches improve accuracy, they struggle in large-scale repositories due to concern mixing, where relevant logic is buried in large functions, and concern scattering, where related logic is dispersed across files.
@@ 1. Introduction
More broadly, the challenges of issue localization and the limitations of existing approaches can be summarized as follows. In large-scale code repositories, localization is hindered by the lack of a high-level conceptual view and by two core difficulties: concern mixing and concern scattering. Concern mixing arises when critical logic is buried within large, multi-purpose functions, making it difficult to isolate the relevant code. Concern scattering occurs when semantically related logic is dispersed across multiple files or modules, forcing developers to piece together fragmented functionalities to identify the root cause. A motivating example from the Expensify repository demonstrates that these challenges not only slow down human developers but also restrict the effectiveness of LLM-based approaches such as mini-SWE-agent (Yang et al., 2024a), which frequently mislocalize issues due to inefficient and unguided codebase exploration (see Section 2).
@@ 2. Motivation
In large-scale code repositories, issue localization is often hindered by the absence of a high-level conceptual view, as well as by concern mixing and concern scattering, which make it difficult for existing methods to link abstract issue descriptions to the specific code elements responsible for the problem. Figure 1 illustrates these challenges using a motivating example from a real issue in the Expensify repository. Figure 1(a) shows the issue along with its corresponding fix patch. Figure 1(b) presents a comparison of the analysis processes of mini-SWE-agent (Yang et al., 2024a) under different settings.
@@ 2.1. Challenges in Issue Localization
Concern Mixing Challenge. Within ReportDetailsPage, the code directly related to the issue, the call to leaveChat(), occupies only a tiny fraction of the function, which spans 792 lines and encompasses many other report-related concerns, such as rendering the report header and managing report-specific actions. As a result, the leave-group logic is entangled with numerous unrelated responsibilities. Without explicit separation of concerns, isolating this small but critical piece of functionality becomes difficult, exemplifying the challenge of concern mixing.
@@ 2.2. Approach Comparison and Our Technique
Our Concern-Enhanced Idea. To address these challenges, we propose abstracting conceptual knowledge by first decomposing fine-grained functionalities from original code functions (mitigating concern mixing), and then recomposing them into high-level conceptual concerns (mitigating concern scattering). During issue localization, the most relevant concerns are selected to provide high-level guidance, enabling more efficient and effective analysis of the issue and exploration of the codebase (providing high-level conceptual overview). As illustrated in the bottom process of Figure 1(b), once the constructed concerns are provided, mini-SWE-agent can correctly identify the entry-point function ReportDetailsPage by leveraging the associated concern.
@@ 4.4. RQ3: Ablation Study
Analysis. Overall, leveraging high-level concerns in RepoLens consistently enhances localization performance across different baselines, yielding clear improvements at both file and function levels. Furthermore, our comparison of concern quality indicates that the contextual information captured by Concerns is both more relevant and of higher quality, which accounts for the observed performance gains. Notably, w/o Con outperforms w/o Exp, suggesting that focusing on term-centric functionalities within functions is more effective than relying solely on general function summaries, as it captures term-specific semantics and helps mitigate concern mixing.
@@ 7. Conclusion
We presented RepoLens, an approach that enhances issue localization by abstracting fine-grained functionalities into high-level conceptual concerns. This addresses challenges such as concern mixing and concern scattering and provides LLMs with structured guidance for efficient and accurate code exploration. Experiments show that RepoLens consistently improves state-of-the-art localization tools across multiple models, while ablation studies and manual evaluation confirm the reliability of the constructed concerns. These results demonstrate the effectiveness of leveraging conceptual knowledge for issue localization in large-scale software systems.

### v2

@@ Abstract.
Issue localization, which identifies faulty code elements such as files or functions, is critical for effective bug fixing. While recent LLM-based and LLM-agent-based approaches improve accuracy, they struggle in large-scale repositories due to concern tangling, where relevant logic is buried in large functions, and concern scattering, where related logic is dispersed across files.
@@ 1. Introduction
More broadly, the challenges of issue localization and the limitations of existing approaches can be summarized as follows. In large-scale code repositories, localization is hindered by the lack of a high-level conceptual view and by two core difficulties: concern tangling and concern scattering. Concern tangling arises when critical logic is buried within large, multi-purpose functions, making it difficult to isolate the relevant code. Concern scattering occurs when semantically related logic is dispersed across multiple files or modules, forcing developers to piece together fragmented functionalities to identify the root cause. A motivating example from the Expensify repository demonstrates that these challenges not only slow down human developers but also restrict the effectiveness of LLM-based approaches such as mini-SWE-agent (Yang et al., 2024a), which frequently mislocalize issues due to inefficient and unguided codebase exploration (see Section 2).
@@ 2. Motivation
In large-scale code repositories, issue localization is often hindered by the absence of a high-level conceptual view, as well as by concern tangling and concern scattering, which make it difficult for existing methods to link abstract issue descriptions to the specific code elements responsible for the problem. Figure 1 illustrates these challenges using a motivating example from a real issue in the Expensify repository. Figure 1(a) shows the issue along with its corresponding fix patch. Figure 1(b) presents a comparison of the analysis processes of mini-SWE-agent (Yang et al., 2024a) under different settings.
@@ 2.1. Challenges in Issue Localization
Concern Tangling Challenge. Within ReportDetailsPage, the code directly related to the issue, the call to leaveChat(), occupies only a tiny fraction of the function, which spans 792 lines and encompasses many other report-related concerns, such as rendering the report header and managing report-specific actions. As a result, the leave-group logic is entangled with numerous unrelated responsibilities. Without explicit separation of concerns, isolating this small but critical piece of functionality becomes difficult, exemplifying the challenge of concern tangling.
@@ 2.2. Approach Comparison and Our Technique
Our Concern-Enhanced Idea. To address these challenges, we propose abstracting conceptual knowledge by first decomposing fine-grained functionalities from original code functions (mitigating concern tangling), and then recomposing them into high-level conceptual concerns (mitigating concern scattering). During issue localization, the most relevant concerns are selected to provide high-level guidance, enabling more efficient and effective analysis of the issue and exploration of the codebase (providing high-level conceptual overview). As illustrated in the bottom process of Figure 1(b), once the constructed concerns are provided, mini-SWE-agent can correctly identify the entry-point function ReportDetailsPage by leveraging the associated concern.
@@ 4.4. RQ3: Ablation Study
Analysis. Overall, leveraging high-level concerns in RepoLens consistently enhances localization performance across different baselines, yielding clear improvements at both file and function levels. Furthermore, our comparison of concern quality indicates that the contextual information captured by Concerns is both more relevant and of higher quality, which accounts for the observed performance gains. Notably, w/o Con outperforms w/o Exp, suggesting that focusing on term-centric functionalities within functions is more effective than relying solely on general function summaries, as it captures term-specific semantics and helps mitigate concern tangling.
@@ 7. Conclusion
We presented RepoLens, an approach that enhances issue localization by abstracting fine-grained functionalities into high-level conceptual concerns. This addresses challenges such as concern tangling and concern scattering and provides LLMs with structured guidance for efficient and accurate code exploration. Experiments show that RepoLens consistently improves state-of-the-art localization tools across multiple models, while ablation studies and manual evaluation confirm the reliability of the constructed concerns. These results demonstrate the effectiveness of leveraging conceptual knowledge for issue localization in large-scale software systems.